OpenStudy (anonymous):

if f(x)=12^2-5 and g(x)=x+7 find (f o g)(x)

OpenStudy (imer):

(f o g)(x) is same as f(g(x)), also you missed a variable in f(x)

OpenStudy (anonymous):

i typed exactly what it says

OpenStudy (anonymous):

and i know its the same as f(g(x)) but i dont understand how to get the answer /:

OpenStudy (anonymous):

@alicatforever u have missed variable 'x' in f(x)

OpenStudy (anonymous):

oh your right 0.0 it was 12x^2-5

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

whoops sorry /:

OpenStudy (imer):

@alicatforever: and you said I was wrong :(

OpenStudy (anonymous):

sorry D:

OpenStudy (paki):

@CO_oLBoY nice one mate...

OpenStudy (anonymous):

i thought you meant a number for some reason. and i was like you've seen this question before!? what is this sorcery!?

OpenStudy (anonymous):

now fog(x) = f (x+7) = 12 (x+7)^2 - 5 = 12 (x^2 + 49 + 14x) - 5 = 12x^2 +588 +14x - 5 = 12x^2 + 14x + 583

OpenStudy (anonymous):

oh my gosh thank you!

OpenStudy (anonymous):

welcome :) anytime